OpenStudy (anonymous):

The function f(x) = 16(2)x represents the growth of a bee population every year in a remote swamp. Jennifer wants to manipulate the formula to an equivalent form that calculates two times a year, not just once a year. Which function is correct for Jennifer's purpose, and what is the new growth rate?

OpenStudy (anonymous):

@tkhunny

OpenStudy (tkhunny):

Can't understand the formula. \(f(x) = 16\cdot 2^{x}\), where x is an integer.

OpenStudy (anonymous):

f(x) = 16(2)x; growth rate 200% f(x) = 16(2)2x; growth rate 8% f(x) = 16(1.41)x; growth rate 8% f(x) = 16(1.41)2x; growth rate 41%

OpenStudy (anonymous):

@tkhunny these are my options

OpenStudy (tkhunny):

You didn't answer my question. Formatting and communication are important. 1) Why would the growth rate change? We are writing a function to reproduce the previous values. 2) IF my f(x) is correct, then f(1) = 16*2 = 32 and the growth rate is 100% / year. Which formula does that? 3) \(\sqrt{2} = 1.414213652373...\) or about 1.41. Thus, \(g(x) = 16\cdot\left[\sqrt{2}\right]^{x}\), where x is in HALF YEARS produces a SEMI-ANNUAL growth rate of about 41%. 4) Very strange question.